# Chemotaxis-induced phase separation

### v1

@@ Abstract
Chemotaxis is a prevalent strategy in biology that enables both microbes and single (tissue) cells to move collectively and directionally in response to chemical gradients [1, 2, 3, 4, 5]. This can lead to self-organized aggregation when cells respond to gradients of self-secreted signaling molecules. The basic model describing chemotactic aggregation was introduced by Keller and Segel [6]. Diverse model variants account for different chemotactic responses, crowding effects, and cell proliferation [7]. Without cell proliferation and death the chemotactic aggregates are observed to undergo a coarsening process [7, 8, 9, 10]. In contrast, cell growth and death induce a fixed pattern wavelength leading to intricate patterning [11, 7], and even spatiotemporal chaos [12, 13]. However, a unifying understanding of chemotactic aggregates and their long-time dynamics in general Keller–Segel models is missing. It remains unknown whether a generic mechanism links coarsening, the selection of fixed intrinsic wavelengths, and the spatiotemporally dynamic patterns.
In this letter, we derive the generic behavior of chemotactic aggregates in Keller–Segel models by developing a combined description of the counteracting effects of random (diffusive) and chemotactic (directed) cell movement. Stationary aggregates are thereby constructed in phase space based on a reactive area balance representing a generalized Maxwell construction. We then argue that the resulting net cell movement drives the coarsening of collections of well-separated aggregates. (Slow) cell growth and death interrupt and reverse the coarsening process, resulting in both fully nonlinear steady-state patterns with fixed wavelengths (aggregate sizes) and spatiotemporally dynamic aggregates. Our theory is independent of the specific mathematical form of the models and relates the macroscopic dynamics of chemotactic aggregates to (active) phase separation [14, 15, 16] and (mass-conserving) reaction–diffusion systems [17, 18].
@@
$\partial_{t}\rho(\bm{x},t)=D_{\rho}\nabla^{2}\rho-T\nabla\big{[}\chi_{\rho}(\rho)\chi_{c}(c)\rho\nabla c\big{]}+\varepsilon s(\rho).$ (1)
Here, ${T\chi_{\rho}(\rho)\chi_{c}(c)>0}$ represents the strength of chemotaxis with the positive sensitivity functions $\chi_{\rho}$ and $\chi_{c}$. For example, $\chi_{\rho}$ may describe crowding effects that suppress chemotaxis, while $\chi_{c}$ may capture nonlinear kinetics between the chemoattractant and its receptor in the cells [7]. Cell growth and death are described by $s(\rho)$ with a dimensionless strength $\varepsilon$. The dynamics of the chemoattractant, produced by the cells and subject to degradation, is given by ${\partial_{t}c(\bm{x},t)=D_{c}\nabla^{2}c+f(\rho,c)}$. To ensure a net chemoattractant production by the cells, we assume ${\partial_{\rho}f>0}$.
@@
$\partial_{t}\rho(\bm{x},t)=T\nabla\big{[}\chi_{\rho}\rho\nabla\eta\big{]}+\varepsilon s(\rho)\,,$ (2)
in terms of the mass-redistribution potential (cf. Refs. [19, 20, 21])
$\eta(\bm{x},t)=\frac{D_{\rho}}{T}\int\mathrm{d}\rho\,\frac{1}{\chi_{\rho}\rho}-\int\mathrm{d}c\,\chi_{c}\,.$ (3)
Comparing the fluxes in Eq. (2) with the chemotaxis term in Eq. (1), it is evident that $\eta(\mathbf{x},t)$ acts as an effective chemorepellent. Random cell motion, appearing as a separate diffusion term in Eq. (1), is included in the population dynamics, Eq. (2), as one of two terms in the effective chemotaxis field $\eta$. Thereby, diffusion is treated as a modification to the chemotactic response of the cells, and the effective field $\eta$ integrates the counteracting effects of diffusion and chemoattraction.
@@
Mathematically, the mass-redistribution potential $\eta$ resembles the chemical potential in binary phase separation with the mobility $\chi_{\rho}\rho$ [cf. Eq. (2) for ${\varepsilon=0}$ with Model B dynamics] [22, 23]. However, the mass-redistribution potential is in general not the derivative of a free energy functional but follows its own dynamic equation, as detailed in the Supplemental Material [24]. For particular model choices one can find a relation to a mathematical variational structure [25, 26]. For numerical analysis, we specifically consider the minimal Keller–Segel (mKS) model with ${\chi_{\rho}=\chi_{c}=1}$ and a linear chemoattractant production and decay term ${f=\rho-c}$ [27]. Then, one has ${\eta=\frac{D_{\rho}}{T}\log\rho-c}$. We start our analysis with the mass-conserving dynamics where ${\varepsilon=0}$.
@@
Figure 1: (a,b) Illustration of the lateral instability. The mass-redistribution potential (orange) induces redistribution (orange arrows) of the cell density (blue) that counteracts (a) or amplifies (b) the initial perturbation (black arrows). (c) A $\mathsf{\Lambda}$-shaped nullcline leads to stationary peak patterns. The low-density intersection between the nullcline (black) and the flux-balance subspace ${\eta=\eta_{\mathrm{stat}}}$ (orange) determines the low-density plateau $\rho_{-}$ (pattern minimum ${\check{\rho}\gtrsim\rho_{-}}$). The value $\eta_{\mathrm{stat}}$ is determined by the reactive area balance (balance of the red-shaded areas). (d) A $N$-shaped nullcline gives a third high-density NC-FBS intersection determining a high-density plateau $\rho_{+}$ (pattern maximum ${\hat{\rho}\lesssim\rho_{+}}$). The resulting pattern is mesa-shaped.
Lateral instability. — A uniform cell density is unstable if chemotactic drift dominates over diffusive spreading. This requires the mass-redistribution potential $\eta$ to decrease where $\rho$ increases, such that gradients in $\eta$ guide motion towards increased cell densities [Eq. (2); cf. orange arrows in Fig. 1(a,b)]. For long-wavelength modulations, mass redistribution is slow compared to local reactions $f$ causing $\eta$ to quickly relax locally to its quasi-steady-state (QSS) $\eta^{*}(\rho)$, given by the nullcline (NC) ${f(\rho,c(\rho,\eta^{*}))=0}$ [insets in Fig. 1(a,b)]. Thus, chemotaxis-induced accumulation destabilizes uniform densities (lateral instability) if
$\partial_{\rho}\eta^{*}(\rho)<0\,.$ (4)
@@
This instability is based on the self-amplifying redistribution of cell mass, and resemble the mass-redistribution instability found in mass-conserving reaction–diffusion systems as well as spinodal decomposition [20, 18]. Because this is a long-wavelength instability, the nullcline-slope criterion, Eq. (4), is exact beyond the QSS approximation; see Ref. [28].
@@
Stationary aggregate profiles. — The stationary, one-dimensional pattern profile $[\tilde{\rho}(x),\tilde{c}(x)]$ can be determined in the $(\rho,\eta)$-phase plane by generalizing local equilibria theory [29, 20]. In stationary state, the diffusive and chemotactic fluxes must balance, i.e., the total flux ${\sim\partial_{x}\eta}$ [cf. Eq. (2)] must vanish everywhere on a domain with no-flux or periodic boundary conditions. Thus, stationary patterns are constrained to the flux-balance subspace (FBS) ${\tilde{\eta}(x)=\eta_{\mathrm{stat}}}$ [Fig. 1(c,d)]. Given that the densities ${\rho,c>0}$ are bounded from below, the (well-separated) patterns are either mesa-shaped and composed of low- and high-density plateaus connected by interfaces [Fig. 1(d)] or peak patterns. For peaks, the maximum density does not saturate in a plateau [Fig. 1(c)]. These patterns are constructed by noting that the profile curvature $\partial_{x}^{2}\tilde{c}$ vanishes (approximately) in plateau regions and at inflection points of the pattern. The stationary chemoattractant equation implies that those regions correspond to FBS-NC intersections. In particular, the low- and high-density FBS-NC intersections $\rho_{\pm}$ of a N-shaped nullcline are approached by the minimum $\check{\rho}$ and maximum $\hat{\rho}$ pattern densities exponentially in the plateau widths [sharp-interface approximation discussed in Supplemental Material [24]; Fig. 1(d)]. Peak patterns occur if the third FBS-NC intersection $\rho_{+}$ does not exist ($\mathsf{\Lambda}$-shaped nullcline) or lies at higher densities (asymmetric N-shaped nullcline) than the total domain mass $M$ [blue-shaded in Fig. 1(c,d)], fixed by the initial conditions, allows for. For example, the mKS model has a $\mathsf{\Lambda}$-shaped nullcline allowing unconstrained cell-density growth while receptor-binding kinetics [30] or volume-filling effects [31, 32] lead to N-shaped nullclines, directly explaining their density saturation. A precise, mathematical criterion for the transition from peak to mesa patterns is given in Ref. [33].
The FBS position $\eta_{\mathrm{stat}}$ was introduced previously as an integration constant [34]. Here, increasing $\eta$ at fixed density $\rho$ corresponds to decreasing the chemoattractant density $c$ [Eq. (3)]. Therefore, chemoattractant is produced (${f>0}$) above and degraded (${f<0}$) below the nullcline in the $(\rho,\eta)$-phase plane (multistability of the reaction kinetics is discussed in the Supplemental Material [24]). Consequently, the stationary value $\eta_{\mathrm{stat}}$ of a pattern adjusts such that the total production and degradation of the chemoattractant is balanced. This balance of the total turnover is represented (qualitatively) by a balance of areas enclosed between FBS and NC analogously to a Maxwell construction [Fig. 1(c,d); see Supplemental Material [24]]. For example, in the mKS model, the total turnover balance reads ${0=\int_{\check{\rho}}^{\hat{\rho}}\mathrm{d}\rho\,[\eta_{\mathrm{stat}}-\eta^{*}(\rho)]/\rho}$.
@@
Figure 2: Illustration of the mass-competition instability for the competition of two peaks (a) and two plateaus (b; coalescence of the peak with its mirror image). In both cases (a) and (b), the changes in the stationary mass-redistribution potentials $\delta\eta$ of the two elementary patterns with the peak mass (blue-shaded areas) and plateau lengths, lead to self-amplifying mass transport (orange). (c) In a large system, these instabilities induce a coarsening process of the pattern as shown in two simulations with $T\bar{\rho}/D_{\rho}=1.5,3$. The faster mass-competition process [(inset in (d)] dominates the dynamics. (d) During the coarsening process, the average pattern wavelength $\langle\Lambda\rangle(t)$ continuously grows. Numerical simulations at $T\bar{\rho}/D_{\rho}=1.25,1.5,2,3$ (average over 3,1,8 and 2 replicates; light green to dark blue) are compared with the prediction ${\langle\Lambda\rangle(t)\sim\max(\gamma_{\pm})\log(t)}$ with ${\gamma\propto 5/8,3/4,1,1}$ (black lines; see Supplemental Material [24]). Apart from the average density $\bar{\rho}$ specified above, the simulation parameters are $T=5$, $D_{\rho}=0.1$, $D_{c}=1$, with system length $L=1000$.
Coarsening due to mass competition. — To understand the dynamics of patterns with several domains (peaks or mesas), we describe the interaction between two elementary patterns (two half-peak or two half-mesa patterns). As illustrated for peaks in the following, the pattern domains can interact either via competition for mass or coalescence with neighbors (mirror image) [Fig. 2(a,b); reflective boundaries]. Assuming a slow evolution of the pattern after each single elementary pattern relaxed to its (quasi-)stationary profile, the mass-redistribution potential at the single interfaces is given by its stationary value $\eta_{\mathrm{stat}}(M_{1,2})$ (QSS approximation). During competition [Fig. 2(a)], the dependence of $\eta_{\mathrm{stat}}(M_{1,2})$ on the peak masses ${M_{1,2}=\bar{M}\pm\delta M}$ induces a gradient $\eta(x)$ that redistributes mass between the peaks. The QSS value $\eta_{\mathrm{stat}}$ of the elementary pattern also depends on the plateau widths. Thus, during coalescence [Fig. 2(b)] the difference in the plateau widths induces a gradient $\eta(x)$ within the peak that induces peak motion.
The mass and plateau-length dependence of $\eta_{\mathrm{stat}}(M)$ arises because the reactive area balance depends on $\check{\rho},\hat{\rho}$, which vary with the peak mass and plateau lengths [Fig. 2(a,b)]. Associating changes in the low-density plateau lengths $\delta l$ with mass changes ${\delta M\approx\pm(\hat{\rho}-\check{\rho})\delta l}$, one describes both effects in terms of the mass dependence of $\eta_{\mathrm{stat}}$. To separate the effects of the peak height and the low-density plateau length, we denote the mass-induced change of the mass-redistribution potential only due to the change in $\hat{\rho}$ or $\check{\rho}$ by $\partial_{M}^{\pm}\eta_{\mathrm{stat}}$, respectively (see Ref. [28]). Importantly, because the reactive area balance does not depend on the specific mathematical model chosen, it generically implies ${\partial_{M}^{\pm}\eta_{\mathrm{stat}}<0}$ for both peak and mesa patterns; assuming for peak patterns that their height $\hat{\rho}$ grows with the mass $M$. Thus, mass is generically redistributed to the elementary pattern with initially higher mass, inducing a mass-competition instability. For peak competition [Fig. 2(a)], this leads to the collapse of the smaller and growth of the larger peak. For plateau competition [Fig. 2(b)], the shorter plateau vanishes due to the coalescence of the peak with its neighbor (mirror image). The growth rates of these two instabilities ${\sigma^{\pm}\sim-\partial_{M}^{\pm}\eta_{\mathrm{stat}}}$ depend on the strength of the $\eta$ gradient induced by the mass differences, i.e., the magnitude of $\partial_{M}^{\pm}\eta_{\mathrm{stat}}$ (see Supplemental Material [24]). Both competition processes drive the coarsening of patterns with several peaks or mesas until only one aggregate remains. Thus, uninterrupted coarsening is generic in mass-conserving Keller–Segel systems analogously to Ostwald ripening and coalescence in phase-separating systems and mass competition in reaction–diffusion systems [23, 17, 18].
In the mKS model, we find ${\sigma^{\pm}\sim\exp(-\Lambda/\gamma_{\pm})}$, $\sigma^{-}$ agreeing with the mathematical result [10]. Because one has ${\gamma_{-}\gtrless\gamma_{+}}$ for ${\bar{\rho}\gtrless 2D_{\rho}/T}$ [Fig. 2(d), inset], coalescence or competition is faster, respectively, and dominates the coarsening process at large wavelengths, which is verified numerically [Fig. 2(c)]. As ${[\sigma^{\pm}(\langle\Lambda\rangle)]^{-1}}$ is the typical timescale on which peaks vanish via collapse or coalescence at the mean pattern wavelength $\langle\Lambda\rangle$, the time during coarsening scales as ${t^{-1}\sim\sigma^{\pm}(\langle\Lambda\rangle)}$ [35, 36, 18]; see Supplemental Material [24]. The resulting coarsening law ${\langle\Lambda\rangle\sim\max(\gamma_{\pm})\log t}$ describes numerical simulations well, with deviations at the transition from competition- to coalescence-dominated coarsening [Fig. 2(d)].
@@
Figure 3: (a) Illustration of interrupted coarsening due to production and degradation [$s(\rho)$; purple] counteracting the mass-competition instability (orange). (b) Plateau splitting occurs when the plateau density enters the laterally unstable region (green) due to production in the low-density plateau. In mesa patterns, high-density plateaus can split as well due to degradation when the plateau density decreases into the laterally unstable density regime. (c) The stability of patterns of the wavelength $\Lambda$ in the mKS model with a logistic source term ${s(\rho)=\rho(1-\rho)}$ of strength $\varepsilon$. Due to self-amplifying production in the low-density plateau ${s(\rho)\sim\rho}$, the threshold of plateau splitting (green) lies at low wavelengths $\Lambda$ and deforms the threshold of interrupted coarsening at large source strength $\varepsilon$ (solid compared to dashed blue line). The analytic thresholds (lines) well describe the numerical thresholds (circles and squares) obtained by numerical continuation and linear stability analysis (see Supplemental Material [24]). (d) The kymographs show $\log(1+\rho)$ (grayscale, 00 to $1$) for $\varepsilon=0.01,0.1$ simulated on a domain with reflective boundaries and length $L=149,74$ corresponding (approximately) to a multiple of the mean of the numerical thresholds $\Lambda_{\mathrm{stop,split}}$. In panels c and d, the same simulation parameters are used as in Fig. 2(d).

### v2

@@ Abstract
Chemotaxis is a prevalent strategy in biology that enables both microbes and tissue cells to move collectively and directionally in response to chemical gradients [1, 2, 3, 4, 5]. This can lead to self-organized aggregation when cells respond to gradients of self-secreted signaling molecules, described by the foundational Keller–Segel (KS) model [6]. Diverse model variants account for different chemotactic responses, crowding effects, and cell proliferation [7]. Without cell proliferation and death the chemotactic aggregates are observed to undergo a coarsening process [7, 8, 9, 10, 11, 12]. In contrast, cell growth and death induce a fixed pattern wavelength leading to intricate patterning [13, 7] and spatiotemporal chaos [14, 15], as well as an absorbing-state phase transition with distinct scaling regimes [16]. However, a unifying understanding of chemotactic aggregates and their long-time dynamics in general KS models is missing. It remains unknown whether a generic mechanism links coarsening, the selection of fixed intrinsic wavelengths, and the spatiotemporally dynamic patterns.
In this letter, we derive the generic behavior of chemotactic aggregates in KS models by developing a combined description of the counteracting effects of random (diffusive) and chemotactic (directed) cell movement. Stationary aggregates are thereby constructed in phase space based on a reactive area balance representing a generalized Maxwell construction. We then argue that the resulting net cell movement drives the coarsening of collections of well-separated aggregates. (Slow) cell growth and death interrupt and reverse the coarsening process, resulting in both fully nonlinear steady-state patterns with fixed wavelengths (aggregate sizes) and spatiotemporally dynamic aggregates. Our theory is independent of the specific mathematical form of the models and relates the macroscopic dynamics of chemotactic aggregates to (active) phase separation [17, 18, 19] and (mass-conserving) reaction–diffusion systems [20, 21].
@@
$\partial_{t}\rho(\mathbf{x},t)=D_{\rho}\nabla^{2}\rho-T\nabla\big[\chi_{\rho}(\rho)\chi_{c}(c)\rho\nabla c\big]+\varepsilon s(\rho).$ (1)
Here, $T\chi_{\rho}(\rho)\chi_{c}(c)>0$} represents the strength of chemotaxis with the positive sensitivity functions $\chi_{\rho}$ and $\chi_{c}$. For example, $\chi_{\rho}$ may describe crowding effects that suppress chemotaxis, while $\chi_{c}$ may capture nonlinear kinetics of the chemotaxis receptors [7]. Cell growth and death are described by $s(\rho)$ with a dimensionless strength $\varepsilon$. The dynamics of the chemoattractant, produced by the cells and subject to degradation, is given by $\partial_{t}c(\bm{x},t)=D_{c}\nabla^{2}c+f(\rho,c)$}. To model net chemoattractant production by the cells, we assume $\partial_{\rho}f>0$}, biasing $f$ toward chemoattractant production with increasing cell density $\rho$.
@@
$\partial_{t}\rho(\mathbf{x},t)=T\nabla\big[\chi_{\rho}\rho\nabla\eta\big]+\varepsilon s(\rho)\,$, (2)
introducing a mass-redistribution potential (cf. Refs. [22, 23, 24])
$\eta(\rho(\mathbf{x},t),c(\mathbf{x},t)):=\frac{D_{\rho}}{T}\int_{\rho_{0}}^{\rho(\mathbf{x},t)}\!\!\!\!\!\!\mathrm{d}\rho\,\frac{1}{\chi_{\rho}\rho}-\int_{c_{0}}^{c(\mathbf{x},t)}\!\!\!\!\!\!\mathrm{d}c\,\chi_{c}\,.$ (3)
Here, we defined two arbitrary reference densities $\rho_{0}$ and $c_{0}$. The mass-redistribution potential is a function of the local densities $\rho$ and $c$ and thus is a spatially varying field. Comparing the sign of the flux term in Eq. (2) with the sign of the chemotaxis term in Eq. (1), it is evident that $\eta(\mathbf{x},t)$ acts as an effective chemorepellent. The effective field $\eta$ integrates the counteracting effects of diffusion and chemoattraction: it increases with $\rho$, promoting spreading from high to low densities, and decreases with $c$, driving aggregation toward regions of high chemoattractant concentration [Eq. (3)].
@@
Mathematically, the mass-redistribution potential $\eta$ resembles the chemical potential in binary phase separation with the mobility $\chi_{\rho}\,\rho$ [cf. Eq. (2) for $\varepsilon=0$} with Model B] [25, 26]. However, $\eta$ is not the derivative of a free energy functional but follows its own dynamic equation; see [27]. For particular model choices, one can find a relation to a mathematical variational structure [28, 29]. The numerical analysis employs the minimal Keller–Segel (mKS) model with $\chi_{\rho}=\chi_{c}=1$} and a linear chemoattractant production and decay term $f=\rho-c$} [30]. Then, one has $\eta=\frac{D_{\rho}}{T}\log\rho-c$}. We start our analysis with the mass-conserving dynamics where $\varepsilon=0$}.
@@
Figure 1: The mass-redistribution potential (orange) induces redistribution (orange arrows) of the cell density (blue) that counteracts (a) or amplifies (b) initial perturbations (black arrows). (c) $\mathsf{\Lambda}$-shaped NCs $\eta^{*}(\rho)$ lead to stationary peak patterns. The low-density intersection between the NC (black) and the flux-balance subspace $\eta=\eta_{\mathrm{stat}}$} (orange) determines the low-density plateau $\rho_{-}$ (pattern minimum $\check{\rho}\gtrsim\rho_{-}$}). The value $\eta_{\mathrm{stat}}$ is determined by the reactive area balance (balance of the red-shaded areas). (d) NN-shaped NCs give an additional high-density NC-FBS intersection $\rho_{+}$ (pattern maximum $\hat{\rho}\lesssim\rho_{+}$}). The resulting pattern is mesa-shaped.
Lateral instability. — A uniform cell density (homogeneous steady state, hss) is unstable if chemotactic drift dominates over diffusive spreading, amplifying small density perturbations. From Eq. (2), this requires that $\eta$ decreases where $\rho$ increases, driving net flux toward denser regions [orange arrows in Fig. 1(a,b)]. Although $\eta$} increases with $\rho$} at fixed chemoattractant concentration c{c} (diffusive contribution), cell-density perturbations also alter c{c} through production, potentially causing $\eta$} to decrease where $\rho$} increases [chemotactic contribution, second term in Eq. (3)]. For long-wavelength modulations, mass redistribution is slow compared to the local relaxation of c{c}, which quickly adjusts to a quasi-steady state (QSS) $c^{*}(\rho)$} defined by the nullcline $f(\rho,c^{*}(\rho))=0$}. Accordingly, $\eta$} locally relaxes to $\eta^{*}(\rho):=\eta(\rho,c^{*}(\rho))$}, and chemotactic aggregation becomes self-amplifying if the effective slope is negative:
$\partial_{\rho}\eta^{*}(\rho_{\mathrm{hss}})<0\,.$ (4)
@@
This nullcline-slop criterion also holds in the short-wavelength regime (see Appendix A). The instability resembles the mass-redistribution instability found in mass-conserving reaction–diffusion systems as well as spinodal decomposition [23, 21].
@@
Stationary aggregate profiles. — The stationary, one-dimensional pattern profile $\tilde{\rho}(x),\tilde{c}(x)]$ can be constructed in the $(\rho,\eta)$-phase plane by generalizing local equilibria theory [31, 23]. In steady state, the diffusive and chemotactic fluxes must balance, i.e., the total flux $\sim\partial_{x}\eta$} [cf. Eq.(2)] must vanish throughout domains with no-flux or periodic boundary conditions. Thus, stationary patterns are constrained to the flux-balance subspace (FBS) $\tilde{\eta}(x)=\eta_{\mathrm{stat}}$} [Fig.1(c,d); $\eta_{\mathrm{stat}}$ is discussed below]. (Well-separated) aggregates are either mesa-shaped—with low- and high-density plateaus connected by interfaces [Fig.1(d)]—or peaks, where density maxima do not saturate in plateaus [Fig.1(c)]. Plateaus and inflection points, where the curvature $\partial_{x}^{2}\tilde{c}$ (approximately) vanishes, are determined by FBS–NC intersections and serve as landmarks of the pattern. In particular, the outer intersections $\rho_{\pm}$ of a N-shaped NC are approached exponentially by the pattern extrema $\check{\rho}$ and $\hat{\rho}$ as plateau widths increase [sharp-interface approximation in [27]; Fig.1(d)]. Peak patterns occur if the rightmost FBS–NC intersection $\rho_{+}$ does not exist (as in a $\mathsf{\Lambda}$-shaped NC) or lies at higher densities (asymmetric N-shaped NC) than the total domain mass $M$ set by initial conditions allows for [blue-shaded in Fig.1(c,d)]. For example, the mKS model has a $\mathsf{\Lambda}$-shaped NC allowing unrestricted cell-density growth, while receptor-binding kinetics [32] or volume-filling effects [33, 34] lead to N-shaped NCs, directly explaining their density saturation. A mathematical criterion for the transition from peak to mesa patterns is given in Ref. [35].
The reaction NC $\eta^{*}(\rho)$ separates the $(\rho,\eta)$ phase plane into regions of net chemoattractant production (f>0{f>0}) and degradation (f<0{f<0}). Since $\eta$ decreases with $c$ at fixed $\rho$ [Eq. (3)], points above the NC correspond to degradation, and points below to production (multistability is discussed in [27]). The FBS position $\eta_{\mathrm{stat}}$—previously introduced as an integration constant [36]—sets the vertical placement of the pattern in this phase plane, and thus the relative contributions of production and degradation. For the pattern to be stationary, $\eta_{\mathrm{stat}}$ must self-adjust to balance total production and degradation. This turnover balance is captured by a Maxwell-like area rule between FBS and NC [Fig. 1(c,d); see [27]]; for example, in the mKS model: $0=\int_{\check{\rho}}^{\hat{\rho}}\mathrm{d}\rho\,[\eta_{\mathrm{stat}}-\eta^{*}(\rho)]/\rho$}.
@@
Figure 2: Illustration of the mass-competition instability for the competition of two peaks (a) and two plateaus (b; coalescence of the peak with its mirror image) destabilizing the symmetric pattern (dashed lines). In both cases (a) and (b), the changes in the stationary mass-redistribution potentials $\delta\eta$ of the two elementary patterns with the peak mass (blue-shaded areas) and plateau lengths lead to self-amplifying mass transport (orange). The generalized Maxwell area balance (shown for the right elementary patterns in the lower panels) shows that the (quasi-)stationary mass-redistribution potential decreases at the larger peak and shorter plateau (red- and green-shaded areas). (c) In a large system, these instabilities induce a coarsening process of the pattern by competition (examples of peak collapse marked by red arrows) and coalescence (blue arrows) as shown in two kymographs of simulations with $T\bar{\rho}/D_{\rho}=1.5,3$. The faster mass-competition process [cf. inset in (d)] dominates the dynamics. (d) During the coarsening process, the average pattern wavelength $\langle\Lambda\rangle(t)$ continuously grows. Numerical simulations at $T\bar{\rho}/D_{\rho}=1.25,1.5,2,3$ (average over 3, 1, 8, and 2 replicates; light green to dark blue) are compared with the prediction $\langle\Lambda\rangle(t)\sim\max(\gamma_{\pm})\log(t)$} with $\gamma\propto 5/8,3/4,1,1$} (black lines; see [27]). Apart from the average density $\bar{\rho}$ specified above, the simulation parameters are $T=5$, $D_{\rho}=0.1$, $D_{c}=1$, with system length $L=1000$.
Coarsening due to mass competition. — In large systems, the KS systems form multiple (quasi-stationary) aggregates. Intriguingly, a multitude of numerical and mathematical work on specific models showed that these multiple aggregates are not stable but undergo a coarsening process, paralleling Ostwald ripening and coalescence of droplets in phase-separating mixtures [10, 12, 37, 11, 38, 39, 8, 9]. Our framework based on the reactive area balance identifies the model-independent mechanism that explains uninterrupted coarsening in mass-conserving KS systems.
Coarsening proceeds via competition between and coalescence of neighboring aggregates, exemplified for peak patterns in Fig. 2(a,b). Our analysis (see Appendix B) shows that these processes are caused by mass competition: the self-amplifying mass transport along gradients in $\eta$, either between aggregates or from one side of the aggregate toward the other. Mass competition arises because the reactive area balance generically implies that $\eta_{\mathrm{stat}}$ decreases both with the mass $M=\int\mathrm{d}x(\rho-\rho_{-})$} (height) of half-mesas and -peaks (elementary patterns)—at fixed length of the low-density plateau [Fig. 2(a)], and also with decreasing length of the low-density plateau—at fixed mass [Fig. 2(b)]. Within a QSS approximation, $\eta$ thus decreases during mass competition at the elementary pattern with the already larger peak mass [Fig. 2(a)] or shorter low-density plateau [Fig. 2(b)], resulting in additional mass transport toward these and an amplification of any initial asymmetry. We quantify the growth rates $\sigma_{\pm}>0$} of this mass-competition instability in the competition and coalescence scenario, respectively, in Appendix B and extend the analysis beyond the QSS approximation in Ref. [40]. This instability makes uninterrupted coarsening a generic feature of mass-conserving Keller–-Segel systems and allows its quantification [see Fig. 2(c,d) and Appendix C]. It unifies previous results for specific systems [38, 39, 8, 41, 42, 10, 12] and likewise accounts for phase-separation-like dynamics in systems where chemotactic drift arises from quorum-sensing-regulated motility [43, 44] or from cross-diffusion couplings [24]. Thus, our framework establishes chemotaxis-induced phase separation as part of the broader class of coarsening phenomena in nonequilibrium pattern-forming systems.
@@
Figure 3: (a) Illustration of interrupted coarsening due to production and degradation $s(\rho)$; purple] counteracting the mass-competition instability (orange) which destabilizes the symmetric peak position against shifts $\delta x$. (b) Plateau splitting occurs when the plateau density enters the laterally unstable region (green) due to production in the low-density plateau. In mesa patterns, high-density plateaus can split as well due to degradation when the plateau density decreases into the laterally unstable density regime. (c) The stability of patterns with wavelength $\Lambda$ in the mKS model with a logistic source term $s(\rho)=\rho(1-\rho)$} of strength $\varepsilon$. The threshold of plateau splitting (green) lies at low wavelengths $\Lambda$ and deforms the threshold of interrupted coarsening at large source strength $\varepsilon$ (solid compared to dashed blue line). The analytic thresholds (lines) well describe the numerical thresholds (circles and squares) obtained by numerical continuation and linear stability analysis (see [27]). (d) The kymographs show $\log(1+\rho)$ (grayscale, 0 to 11) for $\varepsilon=0.01,0.1$ simulated on a domain with reflective boundaries and length $L=149,74$ corresponding (approximately) to a multiple of the mean of the numerical thresholds $\Lambda_{\mathrm{stop,split}}$. The same simulation parameters are used as in Fig. 2(d).